The Metropolitan City Hospital is unionized. In 2011, nurses receired an average of $45,000, The hospital administrator is considering how the con- tract with nurses should be changed for 2012. In turn, the current practice of charging nursing costs to each department might also be changed. Each departmentis accountable for its financial performance. Revenues and xpenses are allocated to departments. Consider the expenses of the obstetrics department in 2011 Variable expenses (based on 2011 patient-days) are Meals 510,000 Laboratory 900,000 Pharmacy 800,000 Maintenance 150,000 Other530,000 Total $3,150,000 Fixed expenses (based on number of beds) are Rent General administrative Janitorial $3,000,000 services 2,100,000 200,000 150,000 450,000 S5,900,000 Other Total Narses are assigned to departments on the basis of annnal patient-days, as Volume Level Number of Patient-Days 10,000 12,000 12,001-16,000 Nurses 30 35 Total patient-days are the number of patients multiplied by the mumber of days they are hospitalized. Each department is charged for the salaries of the murses assigned to it. department had a capacity of 60 beds, billed average of $800 per day for each During 2011 the o insurance and government agencies an and had revenues of $12 million. 1. Compute the 2011 volume of activity in patient-days. 2. Compute the 2011 patient-days that would have been 3. Compute the 2011 patient-days that would have been necessary for the obstetrics 4. Suppose Obstetrics must pay $200 per patient-day for nursing services. This plan necessary for the obstetrics department to recoup all fixed department to break even, including nurses salaries as a fixed cost would replace the two-level fixed-cost system employed in 2011 ercept nursing expenses. break-even point in patient-days would have been in 2011 under this
Answer #1

1.Capacity = 60 beds

Revenues = $12,000,000

Billing per patient per day = $800

Patient Days = Revenue/Bill per patient per day

= 12,000,000/800

= 15,000

2.Total Fixed Expenses except nursing expenses = $5,900,000

Bill per patient per day = $800

Variable Cost per patient per Day = $3,150,000/15,000

= $210

Contribution Margin per patient day = $590

Required patient days = 5,900,000/590

= 10,000

3.Including nursing cost of 10,000-12,000 volume level, nursing cost = 30*45,000 = $1,350,000

Other Fixed Costs = $5,900,000

Total Fixed Costs = $7,250,000

Bill per patient per day = $800

Variable Cost per patient per Day = $3,150,000/15,000

= $210

Contribution Margin per patient day = $590

Required Patient Days = 7,250,000/590

= 12,288.14

Since the volume level is more than 12,000, nursing cost will be higher

Break even point will be calculated as:

Including nursing cost of 12,001-16,000 volume level, nursing cost = 35*45,000 = $1,575,000

Other Fixed Costs = $5,900,000

Total Fixed Costs = $7,475,000

Bill per patient per day = $800

Variable Cost per patient per Day = $3,150,000/15,000

= $210

Contribution Margin per patient day = $590

Required Patient Days = 7,475,000/590

= 12,669.50

4.Revenue per patient day = $800

Variable Cost per patient day = $210

Nursing cost per patient day = $200

Contribution Margin per patient day = $390

Total Fixed Costs = $5,900,000

Break-even point = 5,900,000/390

= 15,128.21 patient days